OpenStudy (emmyma):

I will give a medal and become a fan! Please help :) If y varies directly with x and y = -7 when x = -1, find x when y = -84. Thanks so much! :)

OpenStudy (cggurumanjunath):

Y=KX

OpenStudy (cggurumanjunath):

-7=K(-1)

OpenStudy (cggurumanjunath):

K=7

OpenStudy (cggurumanjunath):

Y=KX -84=7*X X=-84/7

OpenStudy (emmyma):

Wow, thank you so much for explaining that so well! :)

OpenStudy (cggurumanjunath):

Anytime. :)